OpenStudy (anonymous):

Oi! Two bags are full of quarters, of which some are American and some are Canadian. The first bag has 169 quarters, of which 13 are Canadian. The second bag has 21 Canadian quarters. If both bags have the same fraction of Canadian quarters, how many quarters does this other bag have?

OpenStudy (anonymous):

13/169=23/x x=299

OpenStudy (anonymous):

441

OpenStudy (anonymous):

i got 273

OpenStudy (anonymous):

169/13=13 this means that you square the other number (21) which gives you 441

OpenStudy (anonymous):

273 is right

OpenStudy (anonymous):

97 227 109 237 273 One of these is the answer, this is what it shows on the answer sheet

OpenStudy (anonymous):

Yes 273 is the correct i just got it, thank you guys

OpenStudy (anonymous):

wow ! ok, so i need to brush up on my math skills! LOL srry!